# TextCrafter: Accurately Rendering Multiple Texts in Complex Visual Scenes

### v2

@@ 3.4 Text Focus
Region insulation creates distinct visual boundaries between visual texts. To enhance harmony, regions are merged into a single latent variable for denoising, but this may blur smaller texts. To preserve structure, we boost attention scores Hertz et al. (2022) for texts and preceding quotation marks, capping enhancement at 2$\times$ to prevent overamplification. To ensure control, we employ the $\tanh$ function. The $k$ token sequence is denoted as $F=\{f_{1},\dots,f_{k}\}$. The enhancement ratio is:
@@
We use a smaller ratio for single-word cases. Unlike UNet’s cross-attention, MM-DiT employs a full attention matrix with four regions: image-to-image, prompt-to-prompt, prompt-to-image, and image-to-prompt Cai et al. (2024) . We re-weight the image-to-text matrix, which controls image modality and enhances focus on visual text. At step $t$, the operation for each element in $M_{t}$ is expressed as:
@@
Step 1 Step 2 Step 3 Word Accuracy↑ NED↑
$\checkmark$ 0.4422 0.6495
$\checkmark$ 0.6116 0.7750
$\checkmark$ 0.6351 0.8055
$\checkmark$ $\checkmark$ 0.6121 0.7762
$\checkmark$ $\checkmark$ 0.7328 0.8651
$\checkmark$ $\checkmark$ $\checkmark$ 0.7370 0.8679
Table 3: Ablation on CVTG-2K of Instance Fusion  (Step 1), Region Insulation  (Step 2), and Text Focus  (Step 3).
@@ 4.1 Benchmark
Currently, there is a lack of publicly available benchmark datasets specifically designed for complex visual text generation tasks. Therefore, we propose CVTG-2K, a challenging public benchmark dataset for complex visual text. All prompts are generated through the OpenAI’s O1-mini API OpenAI (2024). These prompts encompass a variety of scenes containing complex visual texts, including but not limited to street views, book covers, advertisements, posters, notes, memes, logos, and movie frames. Unlike previous standard datasets synthesized using fixed rules, CVTG-2K ensures the diversity and rationality of the data distribution. We first prompt the O1 model to conceive a scene and then imagine the spossible visual texts that may appear in that scene, utilizing Chain-of-Thought (CoT) Wei et al. (2022) techniques to ensure the quality of the generated prompts. For further details on the dialogue template with O1, refer to the supplementary material. Through rigorous filtering rules and meticulous post-processing, we have generated 2,000 prompts containing complex visual texts. On average, the visual texts in CVTG-2K contain 8.10 words and 39.47 characters, surpassing all previously published visual text benchmark datasets in terms of visual text length. Table 1 presents a comparison of the data in CVTG-2K with currently available benchmark datasets. Furthermore, CVTG-2K is the first benchmark dataset to include the number of multiple visual text regions, with the number of regions ranging from 2 to 5. We have assigned different proportions to these region numbers, approximately 20%, 30%, 30%, and 20%. With its diverse region numbers and visual text lengths, CVTG-2K provides a comprehensive evaluation of model performance in complex visual text generation tasks. Through careful human review and filtering, CVTG-2K ensures that it does not contain any discriminatory or inflammatory content.
@@
To further enhance the challenge of CVTG-2K, we randomly added attributes to each visual text in half of the data. The attributes include size, color, and font. The sizes are categorized as large, medium, and small, and the fonts include regular, bold, italic, and cursive. All attributes are labeled using natural language to ensure that the text encoder can process them without the need for special designs. Each visual text will randomly have one or more of these attributes. This design will promote further research into the stylization and customization of visual texts in future studies. Additionally, we provide more fine-grained information. We decouple the prompts using O1-mini, separating descriptions $D=\{d_{1},d_{2},\dots,d_{n}\}$ containing multiple visual texts and expressing the relationship between each visual text and its position through the most critical words, which is the carrier of the visual text. Through meticulous manual review, we ensured the accuracy of this fine-grained information. CVTG-2K will be publicly released alongside our code, and the dataset will be available for any users adhering to the usage guidelines. For more examples of CVTG-2K, please refer to the supplementary material.
@@ 4.2 Evaluation Metrics and Baselines
The quality of CVTG are evaluated from three metrics:
  1. 1.
    Word Accuracy: The generated image is input into an OCR tool for prediction. Here, we use the SOTA open-source OCR model PPOCR-v4 PaddlePaddle (2023).
  2. 2.
    Normalized Edit Distance (NED): NED Tuo et al. (2024) is a more lenient metric than word accuracy, used to measure character-level similarity.
  3. 3.
    CLIPScore: We introduce CLIPScore, which is used to evaluate the model’s ability to follow the text prompt during the generation process Hessel et al. (2021).
We compare our method on CVTG-2K with several state-of-the-art models for visual text generation, including Stable Diffusion 3.5 Large Esser et al. (2024), FLUX.1 dev, AnyText Tuo et al. (2024) and TextDiffuser2 Chen et al. (2025). To highlight the difference between CVTG and multi-instance generation tasks, we also compare with methods like RAG-Diffusion Chen et al. (2024b) and 3DIS Zhou et al. (2025). All experiments use official or author-provided codes and configurations for optimal baseline performance. The parameter $\lambda$ described in Section 3.2 uses 0.4, and $r=5$ from Section 3.3. Notably, TextCrafter is built on FLUX.1 dev and requires no training or fine-tuning, enabling all experiments to run on a single A6000 GPU.
@@ 4.3 Quantitative Results
As shown in Table 2, quantitative experiments on the CVTG-2K demonstrate that TextCrafter outperforms competing methods in both OCR accuracy (Word Accuracy and NED) and prompt-following ability (CLIPScore). Specifically, TextCrafter improves OCR accuracy by over 45% compared to FLUX. While general image generation models like Stable Diffusion 3.5 and FLUX perform well in simple visual text scenarios, their performance significantly degrades as the number of visual text regions increases. Methods like AnyText and TextDiffuser-2, trained on rule-based data and focusing on single-region generation, fail to generalize well to complex visual text tasks. AnyText performs better when the mask covers a larger portion of the image, leading to a slight improvement in accuracy as the number of regions increases. However, its overall performance remains 55% behind that of TextCrafter. Multi-instance control methods such as RAG-Diffusion and 3DIS, although capable of rendering complex scenes, struggle with the specific requirements of visual text generation, resulting in poor performance as complexity rises. Detailed quantitative results are provided in the supplementary material. Overall, the experiments confirm that TextCrafter excels in robustness, maintaining high accuracy and prompt-following performance even in complex visual text scenarios.
@@ 4.4 Qualitative Results
Figure 5 presents visual results comparing TextCrafter with several state-of-the-art text-to-image generation models, including SD3.5, FLUX, AnyText, TextDiffuser-2, RAG-Diffusion and 3DIS, on the CVTG-2K. While SD3.5 and FLUX produce high-quality images and can generate some visual texts, they struggle with omissions and confusion as the number of regions increases. AnyText performs suboptimally when the text consists of two or more words. TextDiffuser sacrifices background information, leading to a lower CLIPScore. RAG-Diffusion suffers from merging issues during the refinement phase. 3DIS weakens visual text information due to its layout-to-depth conversion. In contrast, TextCrafter maintains high image harmony and accurately renders multiple visual texts, adhering to prompt specifications while avoiding attribute confusion. More qualitative results are available in the supplementary material.
@@ 4.5 Ablation Study
Table 3 presents quantitative ablation results of each step individually or in combinations.
@@
Disabling instance fusion has a small impact on overall metrics, as it primarily establishes the correlation between visual text and its location, which doesn’t directly affect accuracy. However, without instance fusion, visual text often appears in incorrect locations. For instance, in the first column of the Figure 6, disabling instance fusion causes the text “Hot Dogs” to appear in the wrong area. Enabling instance fusion also suppresses hallucinated text, ensuring correct placement and preventing irrelevant text from appearing, as shown in the second column of the Figure 6.
@@
Enabling region insulation improves performance across all metrics by decoupling complex visual text, reducing CVTG complexity and maximizing the pre-trained model’s ability. Region insulation achieves over 60% word accuracy on its own and reduces interference between texts. In Figure 6, region insulation prevents interference, improving visual text clarity and accuracy, as seen in the corrected rendering of “Arrival” in the fourth column.
@@
Text focus has the most significant impact on improving metrics, especially for small-size visual text. It enhances both clarity and accuracy, as seen in the clear generation of “Next Stop” on the train in the fifth column in Figure 6, and “VIP Access” in the last column. When enabled alone, text focus achieves 63.51% word accuracy but is about 10% lower than the optimal performance, indicating that it requires support from instance fusion and region insulation for stability, particularly in complex visual text scenarios.
@@ 5 Conclusion
This paper delves into the complex visual text generation (CVTG) task, which involves generating textual content across multiple regions of an image. We propose TextCrafter, a novel training-free framework based on DiT framework, employing step-by-step refinement from prompts to clauses, and finally to specific visual text for precise generation. Key innovations in TextCrafter include instance fusion to link text with its spatial carrier, the region insulation strategy to reduce complexity and prevent text loss, and a text focus approach for accurate small-text rendering. We also construct a new CVTG-2K benchmark dataset for the CVTG task. Extensive experiments demonstrate TextCrafter’s clear advantages against the existing competitors.
Ethical Statement
Any users of our model are prohibited from using it to create false or malicious information.

### v3

@@ 3.4 Text Focus
Region insulation creates distinct visual boundaries between visual texts. To enhance harmony, regions are merged into a single latent variable for denoising, but this may blur smaller texts. To preserve structure, we boost attention scores [11, 15] for texts and preceding quotation marks, capping enhancement at 2$\times$ to prevent overamplification. To ensure control, we employ the $\tanh$ function. The $k$ token sequence is denoted as $F=\{f_{1},\dots,f_{k}\}$. The enhancement ratio is:
@@
We use a smaller ratio for single-word cases. Unlike UNet’s cross-attention, MM-DiT employs a full attention matrix with four regions: image-to-image, prompt-to-prompt, prompt-to-image, and image-to-prompt [2] . We re-weight the image-to-text matrix, which controls image modality and enhances focus on visual text. At step $t$, the operation for each element in $M_{t}$ is expressed as:
@@
Table 3: Ablation studies on CVTG-2K dataset, including Instance Fusion (Step 1), Region Insulation (Step 2), Text Focus (Step 3), and Layout methods.
Steps Layout Method
Metrics S1 S2 S3 S1+S2 S2+S3 All Random Fixed Optimizer
Word Accuracy↑ 0.4422 0.6116 0.6351 0.6121 0.7328 0.7370 0.6900 0.6994 0.7370
NED↑ 0.6495 0.7750 0.8055 0.7762 0.8651 0.8679 0.8374 0.8475 0.8679
@@ 4.1 Benchmark
Currently, there is a lack of publicly available benchmark datasets specifically designed for complex visual text generation tasks. Therefore, we propose CVTG-2K, a challenging public benchmark dataset for complex visual text. All prompts are generated through the OpenAI’s O1-mini API [25]. These prompts encompass a variety of scenes containing complex visual texts, including but not limited to street views, book covers, advertisements, posters, notes, memes, logos, and movie frames. Unlike previous standard datasets synthesized using fixed rules, CVTG-2K ensures the diversity and rationality of the data distribution. We first prompt the O1 model to conceive a scene and then imagine the spossible visual texts that may appear in that scene, utilizing Chain-of-Thought (CoT) [33] techniques to ensure the quality of the generated prompts. For further details on the dialogue template with O1, refer to the supplementary material. Through rigorous filtering rules and meticulous post-processing, we have generated 2,000 prompts containing complex visual texts. On average, the visual texts in CVTG-2K contain 8.10 words and 39.47 characters, surpassing all previously published visual text benchmark datasets in terms of visual text length. Table 1 presents a comparison of the data in CVTG-2K with currently available benchmark datasets. Furthermore, CVTG-2K is the first benchmark dataset to include the number of multiple visual text regions, with the number of regions ranging from 2 to 5. We have assigned different proportions to these region numbers, approximately 20%, 30%, 30%, and 20%. With its diverse region numbers and visual text lengths, CVTG-2K provides a comprehensive evaluation of model performance in complex visual text generation tasks. Through careful human review and filtering, CVTG-2K ensures that it does not contain any discriminatory or inflammatory content.
@@
To further enhance the challenge of CVTG-2K, we randomly added attributes to each visual text in half of the data. The attributes include size, color, and font. The sizes are categorized as large, medium, and small, and the fonts include regular, bold, italic, and cursive. All attributes are labeled using natural language to ensure that the text encoder can process them without the need for special designs. Each visual text will randomly have one or more of these attributes. This design will promote further research into the stylization and customization of visual texts in future studies. Additionally, we provide more fine-grained information. We decouple the prompts using O1-mini, separating descriptions $D=\{d_{1},d_{2},\dots,d_{n}\}$ containing multiple visual texts and expressing the relationship between each visual text and its position through the most critical words, which is the carrier of the visual text. Through meticulous manual review, we ensured the accuracy of this fine-grained information. CVTG-2K has been publicly released alongside our code, and the dataset is available for any users adhering to the usage guidelines. For more examples of CVTG-2K, please refer to the supplementary material.
@@ 4.2 Evaluation Metrics and Baselines
The quality of CVTG are evaluated from five metrics: Word Accuracy and Normalized Edit Distance (NED) [32] are employed to evaluate the accuracy of text rendering. CLIPScore [12], VQAScore [22], and Aesthetics [20] are utilized to assess the quality of the generated images. For detailed configurations of the metrics calculation, please refer to the supplementary materials.
We compare our method on CVTG-2K with several state-of-the-art Text-to-Image models including Stable Diffusion 3.5 Large [9], FLUX.1 dev. To highlight the difference between CVTG and multi-instance generation tasks, we also compare with methods like 3DIS [45]. To highlight the distinction between CVTG and multi-line text generation tasks, we conducted comparative analyses of methods capable of generating multi-line text, such as AnyText [32] and TextDiffuser2 [5]. The parameter $\lambda$ described in Section 3.2 uses 0.4, and $r=5$ from Section 3.3. Notably, TextCrafter requires no training or fine-tuning and is compatible with different DiT backbone networks. We conducted experiments using both FLUX.1 dev and Stable Diffusion 3.5 Large as backbone networks. All experiments can be completed on a single A6000 GPU. For specific data on memory utilization and computational runtime, please refer to the supplementary materials.
@@ 4.3 Quantitative Results
As demonstrated in Table 2, TextCrafter surpasses competing methods in both text accuracy and image quality on CVTG-2K, with TextCrafter(FLUX) improving word accuracy by over 45% relative to FLUX. While Stable Diffusion 3.5 and FLUX perform adequately in simple scenarios, their efficacy diminishes significantly with increased textual complexity. AnyText and TextDiffuser-2, despite being trained on rule-based data, fail to generalize to multi-region tasks, with AnyText performing 55% below TextCrafter even with improved mask coverage. Similarly, 3DIS struggles with visual text generation despite its multi-instance capabilities. As a training-free approach, TextCrafter maintains superior aesthetic scores compared to specialized alternatives while balancing textual accuracy with visual coherence, demonstrating exceptional robustness across complexity levels. Comprehensive results are available in the supplementary material.
@@ 4.4 Qualitative Results
Figure 6 illustrates comparative visual results between TextCrafter and state-of-the-art models (SD3.5, FLUX, AnyText, TextDiffuser-2, and 3DIS) on CVTG-2K. While SD3.5 and FLUX generate visually appealing images, they exhibit deficiencies in text rendering as regional complexity increases. AnyText demonstrates suboptimal performance with multi-word text instances, TextDiffuser-2 compromises background fidelity (resulting in diminished CLIPScore), and 3DIS deteriorates textual information during layout-to-depth conversion. Conversely, TextCrafter achieves superior visual harmony while accurately rendering multiple text regions without attribute confusion or prompt deviation. Additional qualitative evaluations are presented in the supplementary material.
@@ 4.5 Ablation Study
Table 3 presents quantitative ablation results for individual and combined components.
@@
Disabling instance fusion minimally impacts overall metrics but is vital for correct visual text localization; its absence often leads to misplaced text (e.g., "Hot Dogs," Figure 7, first column). Enabling it also suppresses hallucinated text, ensuring accurate placement (Figure 7, second column).
@@
Enabling region insulation improves all metrics by decoupling complex visual text, thereby reducing CVTG complexity and maximizing pre-trained model utility. It achieves over 60% word accuracy independently and reduces inter-text interference, enhancing clarity (e.g., "Arrival," Figure 7, fourth column).
@@
Text focus offers the most significant metric improvement, especially for small visual text, by enhancing clarity and accuracy (e.g., "Next Stop" and "VIP Access," Figure 7, fifth and last columns). Alone, it yields 63.51% word accuracy, roughly 10% below optimal, indicating a need for instance fusion and region insulation for stability in complex scenarios.
Layout Optimizer Effectiveness.
The proposed Manhattan distance-based layout optimization significantly influences final generation by creating non-overlapping bounding boxes aligned with pretrained models’ spatial preferences. Table 3 quantitatively demonstrates its superiority over alternatives like Random or Fixed Layouts.
@@ 5 Conclusion
This paper addresses complex visual text generation (CVTG) by proposing TextCrafter, a novel training-free framework built on the DiT framework. It employs step-by-step refinement from prompts to clauses and finally to specific visual text. Key innovations include Instance Fusion for text-spatial carrier linkage, Region Insulation to reduce complexity and text loss, and a Text Focus approach for accurate small-text rendering. A new CVTG-2K benchmark dataset is also introduced. Extensive experiments demonstrate TextCrafter’s significant advantages over existing methods.
Limitations
TextCrafter has two main limitations: (1) degraded performance with conflicting requirements (e.g., large text in small areas), where balancing spatial constraints and clarity is challenging; and (2) as a training-free method, its text enhancement can introduce artifacts that affect overall image aesthetics, especially in visually dense scenes. Future work will aim to develop mechanisms for high-fidelity text rendering while maintaining the aesthetic integrity of the image.